OpenStudy (anonymous):

integrate dm/dt = 3m for m=5 and t=1?

OpenStudy (anonymous):

the answer is m=5e^(3t-3) but i dont know why

OpenStudy (anonymous):

i solved the equation and got m=ce^3t but i dont know if that's correct

OpenStudy (anonymous):

Yes it is. Now solve for C using the initial condition.

OpenStudy (anonymous):

ok so i plug it in and i get 5=ce^(3*1)

OpenStudy (anonymous):

but where did the 3t-3 come from?

OpenStudy (anonymous):

and the 5

OpenStudy (anonymous):

It's just the way its written 5e^(3t-3)= 5e^(3t)*e^(-3). In your answer the e^(-3) is enclosed in your C. 5=Ce^(3*1) so C=5/e^3

OpenStudy (anonymous):

oh i see it now thanks!!